The daily demand for beef can be modeled by 40.007 D(0) = - 1 + 0.03320.354p million pounds where the price for beef is p dolIs this clearer?
Answer #1

Supply of beef Supply of 5 0 .395 SLP) = 7+53.98 ēron ) p=4.50 51 - 1453.98 0.395x4.50 51 t 53.98 € *7775 51 = 14 153.98)*(0-Demand for beef 40-007 30.3548 OLP) = +0.033e p=4.50 40.007 1+0.033 € 2 0.354x4.5 yo.007 1 +0.033 e!:593 40.007 1+ 0.033)x (46 market equilibrium will occur when Demand = Supply O(P) = S(p) 51 - 1+53.98 ē 0.3958 40.007 1+ 0.033 0.3548 = 1 40.007 1+53- 0-3950 53.98 -0.276 +0.04 2042 +0.042042 20-7498 de error method trail p = 9:25] DLP= 5(P)- D(0) 40-007 H+ 0.033€0.354X69.2